the length of a rectangle is 6 more than the width. the length of the diagonal is 30 meters. find the dimension of the rectangle
Answer by LinnW(1048)

the length of a rectangle is 6 more than the width. the length of the diagonal is 30 meters. find the dimension of the rectangle
(w)^2 + (w+6)^2 = 30^2
w^2 + ( w^2 + 12w + 36 ) = 900
2w^2 + 12w + 36 = 900
add -900 to each side
2w^2 + 12w -864 = 0
divide each side by 2
w^2 + 6w -432 = 0
(w-18)(w+24)=0
So w = 18 or w = -24
Since we need a positive width, w = 18
The length is 6 more or 24
Check (w)^2 + (w+6)^2 = 30^2
(18)^2 + (24)^2 = 30^2
This checks out.